# Quantum Langevin Dynamics for Optimization

### v1

@@ Abstract
We initiate the study of utilizing Quantum Langevin Dynamics (QLD) to solve optimization problems, particularly those non-convex objective functions that present substantial obstacles for traditional gradient descent algorithms. Specifically, we examine the dynamics of a system coupled with an infinite heat bath. This interaction induces both random quantum noise and a deterministic damping effect to the system, which nudge the system towards a steady state that hovers near the global minimum of objective functions. We theoretically prove the convergence of QLD in convex landscapes, demonstrating that the average energy of the system can approach zero in the low temperature limit with an exponential decay rate correlated with the evolution time. Numerically, we first show the energy dissipation capability of QLD by retracing its origins to spontaneous emission. Furthermore, we conduct detailed discussion of the impact of each parameter. Finally, based on the observations when comparing QLD with classical Fokker-Plank-Smoluchowski equation, we propose a time-dependent QLD by making temperature and $\hbar$ time-dependent parameters, which can be theoretically proven to converge better than the time-independent case and also outperforms a series of state-of-the-art quantum and classical optimization algorithms in many non-convex landscapes.
preprint: APS/123-QED
I Introduction
@@
$H_{total}$ $=H+H_{B}$
$=H+\sum_{n}\left(\frac{p_{n}^{2}}{2m_{n}}+\frac{k_{n}}{2}(q_{n}-X)^{2}\right),$ (2)
@@
where $H=p^{2}/(2m)+V$, $p$ is the momentum operator of the system, $x$ is the position operator, and $\eta=\gamma/2m$ represents the characteristic damping rate with system mass $m$, $k$ is the Boltzmann constant. However, it is well known that this master equation is only applicable in high-temperature scenarios and fails to satisfy the positivity constraint of the density operator [41, 42, 43]. Ultimately, we refer to Ref. [41] and employ Lindblad functional (Eq. (15)) to describe the quantum Langevin dynamics, with a detailed introduction provided in Section II.2.
@@
where $\left\langle{V}\right\rangle_{t}$ denotes the expectation value of $V$ with respect to the density matrix of the system, $V(x^{*})$ is the potential value at the global minimum $x^{*}$ and $f(T)$ is a function with respect to $T$ that satisfies $\lim_{T\to 0}f(T)=0$. Consequently, the average energy of the system can approach zero in the low temperature limit. When $V$ is a quadratic form, the evolution of the system has an analytical solution. One interesting discovery related to the quadratic potential is that the system can reach the ground state of the quadratic potential when the evolution time $t\to+\infty$ and $T\to 0$. Formal description can be found in Section III. However, it is always challenging to give a convergence in the general nonconvex case. Despite our best efforts, we are constrained to confine our theoretical discussions to slightly nonconvex cases. Thus, for other nonconvex landscapes, we resort to meticulous numerical experiments to compensate for the absence of theoretical assurances.
@@
Building upon these observations, we find it imperative to explore the realm of time-dependent QLD. Specifically, we recognize the need for pronounced thermal and tunneling effects initially to facilitate escape from local minima. As the system converges to its final stage, these effects need to be gradually deactivated. Consequently, we present a theoretical analysis for time-dependent QLD and propose a cooling strategy for implementation. The convergence of time-dependent QLD is similar to the time-independent one when $\hbar=\hbar(t)$ decreasing monotonically with respect to time and $\eta=\eta(t)$, which says $\left\langle{V}\right\rangle_{t}-V(x^{*})\leq 2kT+O(e^{-\int_{0}^{t}\eta(t^{\prime})\mathrm{d}t^{\prime}})$. For the case of time-dependent temperature, there exists a cooling strategy $T=T(t)$ that guarantees the convergence. In numerical experiments, time-dependent QLD significantly outperforms its time-independent counterpart and surpasses other algorithms, including QHD, QAA, and SGD, across a diverse array of landscapes. In theory, the query complexity of QLD is $\tilde{O}(t)$, which is also among the top tier of all other algorithms.
@@
The rest of this paper is laid out as follows. In Section II, we introduce our notations and provide details for quantum Langevin equation and explicit description of QLD in Eq. (15). Theoretically, we prove the convergence of QLD in convex case (Theorem 1) and slightly nonconvex case (Theorem 2), refer to Section III. Numerically, we first trace back to spontaneous emission to show the essential cause of energy dissipation in QLD, which verifies our intuition, see Section IV.1. Then we verify our theoretical results of quadratic case in Section IV.3 and discuss the roles of each parameter and their relations with thermal and tunneling effects in Section IV.4. In Section V we compare QLD with classical dynamics, which informs the development of time-dependent QLD to fully leverage the advantages inherent in quantum dynamics. In Section V.1, we prove the convergence of time-dependent QLD in convex case. The performance of time-dependent QLD is evaluated in eight representative landscapes, demonstrating superior performance through comparisons with QHD, QAA, and classical algorithms, shown in Section V.6. Finally, we evaluate and compare the query complexity of QLD with other algorithms in Appendix B, demonstrating that QLD is superior to other algorithms in terms of query complexity.
II Preliminaries
@@
Notations. We denote the Hamiltonian of system as $H$, the Hamiltonian of heat bath as $H_{B}$, the Hamiltonian of heat bath and system as $H_{total}=H+H_{B}$. Let $E_{k}$ be the kinetic energy of the system and $V$ be the potential energy of the system, so $H=E_{k}+V$. Given two quantum operators $A$ and $B$, we use $[A,B]=AB-BA$ for denoting the commutation, while $[A,B]_{+}=AB+BA$ for the anti-commutation. We denote the characteristic damping rate of the system as $\eta$ and the temperature of heat bath as $T$. Given a quantum operator $O(t)$, its time-dependent expectation value with respect to density matrix $\rho(t)$ is calculated by $\langle O(t)\rangle_{t}=\mathrm{Tr}(\rho(t)O(t))$.
II.1 Quantum Langevin Dynamics (QLD)
@@
$H_{total}=H+\sum_{n}\left(\frac{p_{n}^{2}}{2m_{n}}+\frac{k_{n}}{2}(q_{n}-X)^{2}\right),$ (7)
@@
[…]
$-f(t-0)x(0)+\xi(t),$ (8)
@@
II.2 Lindblad functional for QLD
Density matrix approach is often used to describe the dynamics of an open quantum system. Based on this point, Caldeira and Leggett proposed a well known master equation named after themselves as the Caldeira-Leggett equation [40, 36] to describe the dynamics of quantum Langevin system. However, the Caldeira-Leggett equation faces some serious probelms, such as the non-positivity of the density matrix and only available for high temperature region [41, 42, 43]. Refs. [41, 46] fixed these problems by introducing a Lindblad equation, which is a more general form of the Caldeira-Leggett equation.
@@ A Lindblad equation can be presented in the standard form:
$\frac{\mathrm{d}\rho}{\mathrm{d}t}=$ $\mathcal{L}(\rho)$
$=$ $-\frac{\mathrm{i}}{\hbar}\left[H,\rho\right]+\sum_{j=1}^{m}\left(2L_{j}\rho L_{j}^{\dagger}-\left[L_{j}^{\dagger}L_{j},\rho\right]_{+}\right).$ (12)
Ref. [41] studied the case where the Hamiltonian $H$ is given by $H=\frac{p^{2}}{2m}+V$ and the unique Lindblad operator $L$ is a linear combination of position operator $x$ and momentum operator $p$,
$L=(\mu x+\mathrm{i}\nu p),~{}~{}L^{\dagger}=(\mu x-\mathrm{i}\nu p),$ (13)
@@
$\frac{\mathrm{d}\rho}{\mathrm{d}t}+\frac{\mathrm{i}}{\hbar}\left[H^{\prime},\rho\right]=$ $-\mu^{2}[x,[x,\rho]]-2\mathrm{i}\mu\nu\left[x,[p,\rho]_{+}\right]$
$-\nu^{2}[p,[p,\rho]],$ (15)
where $H^{\prime}=H-2\mu\nu\hbar xp$. We also call this equation the Lindblad equation of Quantum Langevin Dynamics, abbreviated as QLD, which is also the name of our algorithm. The coordinate space representation of Eq. (15) [41] is also used in the numerical experiments of QLD, stated here for completeness:
$\frac{\partial\rho\left(x,y,t\right)}{\partial t}+\frac{\mathrm{i}}{\hbar}\left[\widetilde{H}(x)-\widetilde{H}^{*}\left(y\right)\right]\rho\left(x,y,t\right)=$ $-\left(\mu^{2}(T)\left(x-y\right)^{2}+\eta\left(x-y\right)\left(\frac{\partial}{\partial x}-\frac{\partial}{\partial y}\right)\right.$
$\left.-\nu^{2}(T)\hbar^{2}\left(\frac{\partial}{\partial x}+\frac{\partial}{\partial y}\right)^{2}\right)\rho\left(x,y,t\right),$ (16)
@@
III Convergence of QLD
In this section, our primary focus is the time-independent scenario of QLD, implying that all parameters $\eta,\hbar,T,m,\Omega,k$ remain constant throughout evolution. In the convex case, the convergence rate of QLD is $O(e^{-\eta t})$ as demonstrated in Theorem 1. Subsequently, we compute the analytical solution of the quadratic potential, the simplest convex function, in Section III.2. The most remarkable discovery related to the quadratic potential is that the system will reach the ground state of the quadratic potential when the evolution time $t\to+\infty$ and the temperature $T\to 0$. Lastly, we prove that QLD continues to converge to the global minimum even in a specific nonconvex scenario, where the objective function satisfies $x\cdot\nabla V(x)\geq rV(x)\geq 0\text{ for }r\in(0,1)$, at a rate of $O(e^{-\eta rt})$. More details can be referred to Theorem 2 of Section III.3.
III.1 Convergence of QLD in the convex case
Our goal is to demonstrate that the expectation value of the objective function $V(x)$ converges to the global minimum $V(x^{*})$ at a defined rate under the evolution of QLD. This necessitates the consideration of the expectation value of the Hamiltonian $H$, as if we can establish that the expectation value of $H$ diminishes, it follows that the expectation value of $V$ will also decrease. With this understanding, we initially present Lemma 1, which enables us to calculate the time evolution of the expectation value of any observable evolving by standard Lindblad equation. After that, we show some commutation relations used in the calculation referring to Lemma 2. Subsequently, we verify the convergence of the expectation value of $H$ in Proposition 2, and with the assistance of this proposition, we validate the convergence of the expectation value of $V$ in Theorem 1.
@@ III.1 Convergence of QLD in the convex case
$\braket{V}_{t}-V(x^{*})\leq 2kT+O(e^{-\eta t}),$ (17)
if the initial density matrix $\rho(0)$ satisfies $\left\langle{H}\right\rangle_{0}-2kT\geq 0$, where $\langle\cdot\rangle_{t}$ denotes the expectation value with respect to the density matrix $\rho(t)$, i.e., $\langle\cdot\rangle_{t}=\operatorname{Tr}(\cdot\rho(t))$.
@@
$\braket{V}_{t}\leq 2kT+O(e^{-\eta t}).$ (18)
Because $\braket{\frac{p^{2}}{2m}}_{t}+\braket{V}_{t}=\braket{H}_{t}$ and all these three terms are greater than or equal to zero, we have
$\braket{V}_{t}\leq\braket{H}_{t}.$ (19)
By Proposition 2, we get $\braket{V}_{t}\leq\braket{H}_{t}\leq 2kT+O(e^{-\eta t})$. ∎
@@
If the desired accuracy is $\varepsilon$, which means we want $\braket{V}_{t}-V(x^{*})\leq\varepsilon$ when $t\to+\infty$. To achieve this, we only need to set $kT=\varepsilon/2$ and ensure the approximation $\hbar\Omega\ll kT$ still holds.
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\langle O(t)\rangle_{t}=$ $\left\langle{\frac{\mathrm{d}}{\mathrm{d}t}O(t)}\right\rangle_{t}+\frac{\mathrm{i}}{\hbar}\langle[H,O(t)]\rangle_{t}$
$+$ $\sum_{j=1}^{m}\left\langle\left(-L_{j}^{\dagger}\left[L_{j},O(t)\right]+\left[L_{j}^{\dagger},O(t)\right]L_{j}\right)\right\rangle_{t}.$ (20)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\langle O(t)\rangle_{t}$ $=\frac{\mathrm{d}}{\mathrm{d}t}\operatorname{Tr}(O(t)\rho(t))$
$=\operatorname{Tr}\left(\frac{\mathrm{d}O(t)}{\mathrm{d}t}\rho(t)\right)+\operatorname{Tr}\left(O(t)\frac{\mathrm{d}\rho(t)}{\mathrm{d}t}\right)$
$=\left\langle{\frac{\mathrm{d}}{\mathrm{d}t}O(t)}\right\rangle_{t}+\operatorname{Tr}\left(O(t)\frac{\mathrm{d}\rho(t)}{\mathrm{d}t}\right).$ (21)
Let’s focus on the second term $\operatorname{Tr}\left((O(t)\frac{\mathrm{d}\rho(t)}{\mathrm{d}t}\right)$. Inserting Eq. (12) into the second term, we get
$\operatorname{Tr}\left(O(t)\frac{\mathrm{d}\rho(t)}{\mathrm{d}t}\right)$
$=$ $\operatorname{Tr}\left(-\frac{\mathrm{i}}{\hbar}O(t)[H,\rho(t)]\right.$
$\left.+\sum_{j=1}^{m}O(t)\left(2L_{j}\rho(t)L_{j}^{\dagger}-L_{j}^{\dagger}L_{j}\rho(t)-\rho(t)L_{j}^{\dagger}L_{j}\right)\right)$
$=$ $-\frac{\mathrm{i}}{\hbar}\operatorname{Tr}\left(O(t)[H,\rho(t)]\right)$
$+\operatorname{Tr}\left(\sum_{j=1}^{m}O(t)\left(2L_{j}\rho(t)L_{j}^{\dagger}-L_{j}^{\dagger}L_{j}\rho(t)-\rho(t)L_{j}^{\dagger}L_{j}\right)\right),$ (22)
where the first term $\operatorname{Tr}\left(O(t)[H,\rho(t)]\right)$ could be simplified as follows:
$\operatorname{Tr}\left(O(t)[H,\rho(t)]\right)$ $=\operatorname{Tr}\left(OH\rho\right)-\operatorname{Tr}\left(O\rho H\right)$
$=\operatorname{Tr}\left(OH\rho\right)-\operatorname{Tr}\left(HO\rho\right)$
$=\operatorname{Tr}\left([O,H]\rho\right)$

### v2

@@ Abstract
We initiate the study of utilizing Quantum Langevin Dynamics (QLD) to solve optimization problems, particularly those nonconvex objective functions that present substantial obstacles for traditional gradient descent algorithms. Specifically, we examine the dynamics of a system coupled with an infinite heat bath. This interaction induces both random quantum noise and a deterministic damping effect to the system, which nudge the system towards a steady state that hovers near the global minimum of objective functions. We theoretically prove the convergence of QLD in convex landscapes, demonstrating that the average energy of the system can approach zero in the low temperature limit with an exponential decay rate correlated with the evolution time. Numerically, we first show the energy dissipation capability of QLD by retracing its origins to spontaneous emission. Furthermore, we conduct detailed discussion of the impact of each parameter. Finally, based on the observations when comparing QLD with the classical Fokker-Plank-Smoluchowski equation, we propose a time-dependent QLD by setting temperature and $\hbar$ as time-dependent parameters, which can be theoretically proven to converge better than the time-independent case and also outperforms a series of state-of-the-art quantum and classical optimization algorithms in many nonconvex landscapes.
1 Introduction
@@
$H_{\text{total}}=H+H_{B}=H+\sum_{n}\left(\frac{p_{n}^{2}}{2m_{n}}+\frac{k_{n}}{2}(q_{n}-X)^{2}\right),$
@@
where $H=p^{2}/(2m)+V$, $p$ is the momentum operator of the system, $x$ is the position operator, and $\eta=\gamma/2m$ represents the characteristic damping rate with system mass $m$, $k$ is the Boltzmann constant. However, it is well known that this master equation is only applicable in high-temperature scenarios and fails to satisfy the positivity constraint of the density operator [29, 19, 44]. Ultimately, we refer to Ref. [29] and employ Lindblad functional (Eq. (14)) to describe the quantum Langevin dynamics, with a detailed introduction provided in Section 2.2.
@@
where $\left\langle{V}\right\rangle_{t}$ denotes the expectation value of $V$ with respect to the density matrix of the system, $V(x^{*})$ is the potential value at the global minimum $x^{*}$ and $f(T)$ is a function with respect to $T$ that satisfies $\lim_{T\to 0}f(T)=0$. Consequently, the average energy of the system can approach zero in the low temperature limit. When $V$ is a quadratic form, the evolution of the system has an analytical solution. One interesting discovery related to the quadratic potential is that the system can reach the ground state of the quadratic potential when the evolution time $t\to+\infty$ and $T\to 0$. Formal description can be found in Section 3. However, it is always challenging to give a convergence in the general nonconvex case. Despite our best efforts, we are constrained to confine our theoretical discussions to slightly nonconvex cases. Thus, for other nonconvex landscapes, we resort to meticulous numerical experiments to compensate for the absence of theoretical assurances.
@@
Building upon these observations, we find it imperative to explore the realm of time-dependent QLD. Specifically, we recognize the need for pronounced thermal and tunneling effects initially to facilitate escape from local minima. As the system converges to its final stage, these effects need to be gradually deactivated. Consequently, we present a theoretical analysis for time-dependent QLD and propose a cooling strategy for implementation. The convergence of time-dependent QLD is similar to the time-independent one when $\hbar=\hbar(t)$ decreasing monotonically with respect to time and $\eta=\eta(t)$, which says $\left\langle{V}\right\rangle_{t}-V(x^{*})\leq 2kT+O(e^{-\int_{0}^{t}\eta(t^{\prime})\mathrm{d}t^{\prime}})$. For the case of time-dependent temperature, there exists a cooling strategy $T=T(t)$ that guarantees convergence. In numerical experiments, time-dependent QLD significantly outperforms its time-independent counterpart and surpasses other algorithms, including QHD, QAA, and SGD, across a diverse array of landscapes. In theory, the query complexity of QLD is $\tilde{O}(t)$, which is also among the top tier of all other algorithms.
@@
The rest of this paper is laid out as follows. In Section 2, we introduce our notations and provide details for quantum Langevin equation and explicit description of QLD in Eq. (14). Theoretically, we prove the convergence of QLD in convex case (Theorem 1) and slightly nonconvex case (Theorem 3), refer to Section 3. Numerically, we first trace back to spontaneous emission to show the essential cause of energy dissipation in QLD, which verifies our intuition, see Section 4.1. Then we verify our theoretical results of quadratic case in Section 4.3 and discuss the roles of each parameter and their relations with thermal and tunneling effects in Section 4.4. In Section 5 we compare QLD with classical dynamics, which informs the development of time-dependent QLD to fully leverage the advantages inherent in quantum dynamics. In Section 5.1, we prove the convergence of time-dependent QLD in convex case. The performance of time-dependent QLD is evaluated in eight representative landscapes, demonstrating superior performance through comparisons with QHD, QAA, and classical algorithms, shown in Section 5.6. Finally, we evaluate and compare the query complexity of QLD with other algorithms in Appendix B, demonstrating that QLD is superior to other algorithms in terms of query complexity.
2 Preliminaries
@@
Notations. We denote the Hamiltonian of system as $H$, the Hamiltonian of heat bath as $H_{B}$, the Hamiltonian of heat bath and system as $H_{\text{total}}=H+H_{B}$. Let $E_{k}$ be the kinetic energy of the system and $V$ be the potential energy of the system, so $H=E_{k}+V$. Given two quantum operators $A$ and $B$, we use $[A,B]=AB-BA$ for denoting the commutation, while $[A,B]_{+}=AB+BA$ for the anti-commutation. We denote the characteristic damping rate of the system as $\eta$ and the temperature of heat bath as $T$. Given a quantum operator $O(t)$, its time-dependent expectation value with respect to density matrix $\rho(t)$ is calculated by $\langle O(t)\rangle_{t}=\mathrm{Tr}(\rho(t)O(t))$.
2.1 Quantum Langevin Dynamics (QLD)
@@
$H_{\text{total}}=H+\sum_{n}\left(\frac{p_{n}^{2}}{2m_{n}}+\frac{k_{n}}{2}(q_{n}-X)^{2}\right),$ (6)
@@
$m\ddot{x}(t)=-V^{\prime}(x(t))-\int_{0}^{t}f(t-t^{\prime})\dot{x}(t^{\prime})\mathrm{d}t^{\prime}-f(t-0)x(0)+\xi(t),$ (7)
@@
2.2 Lindblad functional for QLD
Density matrix approach is often used to describe the dynamics of an open quantum system. Based on this point, Caldeira and Leggett proposed a well-known master equation named after themselves as the Caldeira-Leggett equation (Eq. (3)) [8, 32] to describe the dynamics of quantum Langevin system. However, the Caldeira-Leggett equation faces some serious probelms, such as the non-positivity of the density matrix and only available for high temperature region [29, 19, 44]. Refs. [29, 58] fixed these problems by introducing a Lindblad equation, which is a more general form of the Caldeira-Leggett equation.
@@ A Lindblad equation can be presented in the standard form:
$\frac{\mathrm{d}\rho}{\mathrm{d}t}=\mathcal{L}(\rho)=-\frac{\mathrm{i}}{\hbar}\left[H,\rho\right]+\sum_{j=1}^{m}\left(2L_{j}\rho L_{j}^{\dagger}-\left[L_{j}^{\dagger}L_{j},\rho\right]_{+}\right),$ (11)
where $L_{j}$ for $j\in[m]$ are linear opeartors known as jump operators. Ref. [29] studied the case where the Hamiltonian $H$ is given by $H=\frac{p^{2}}{2m}+V$ and the unique Lindblad operator $L$ is a linear combination of the position operator $x$ and the momentum operator $p$,
$L=\mu x+\mathrm{i}\nu p,~{}~{}L^{\dagger}=\mu x-\mathrm{i}\nu p,$ (12)
@@
$\frac{\mathrm{d}\rho}{\mathrm{d}t}+\frac{\mathrm{i}}{\hbar}\left[H^{\prime},\rho\right]=-\mu^{2}[x,[x,\rho]]-2\mathrm{i}\mu\nu\left[x,[p,\rho]_{+}\right]-\nu^{2}[p,[p,\rho]],$ (14)
where $H^{\prime}=H-2\mu\nu\hbar xp$. We also call this equation the Lindblad equation of Quantum Langevin Dynamics, abbreviated as QLD, which is also the name of our algorithm. The coordinate space representation of Eq. (14[29] is also used in the numerical experiments of QLD, stated here for completeness:
$\frac{\partial\rho\left(x,y,t\right)}{\partial t}$ $+\frac{\mathrm{i}}{\hbar}\left[\widetilde{H}(x)-\widetilde{H}^{*}\left(y\right)\right]\rho\left(x,y,t\right)=-\left(\mu^{2}(T)\left(x-y\right)^{2}\right.$
$\left.+\eta\left(x-y\right)\left(\frac{\partial}{\partial x}-\frac{\partial}{\partial y}\right)-\nu^{2}(T)\hbar^{2}\left(\frac{\partial}{\partial x}+\frac{\partial}{\partial y}\right)^{2}\right)\rho\left(x,y,t\right),$ (15)
@@
Although Ref. [29] does not elucidate the description of Quantum Langevin Dynamics in high dimension explicitly, its generalization is quite natural. For completeness, we write down the expression of high-dimensional QLD here. Suppose that the system is in the $\mathbb{R}^{d}$ space with ${d}$-dimensional position operator $x=(x_{1},x_{2},\ldots,x_{d})^{\top}$ and momentum operator $p=(p_{1},p_{2},\ldots,p_{d})^{\top}$, intriguing the following commutation relations
$[x_{i},p_{j}]=\mathrm{i}\delta_{ij}\hbar,~{}[x_{i},x_{j}]=0,~{}[p_{i},p_{j}]=0,~{}\text{for }i,j=1,2,\ldots,{d}.$ (16)
The Hamiltonian $H$ of high-dimensional QLD is
$H=\frac{p_{1}^{2}+\cdots+p_{d}^{2}}{2m}+V(x_{1},\ldots,x_{d}),$ (17)
where $V(x_{1},\ldots,x_{K})\colon\mathbb{R}^{d}\to\mathbb{R}$ is the potential energy (and also the objective function we want to optimize). The Lindblad operator of the $j$-th dimension in space $\mathbb{R}^{d}$ is
$L_{j}=\mu x_{j}+\mathrm{i}\nu p_{j},\quad\text{for }j=1,2,\ldots,{d},$ (18)
where the parameters $\mu$ and $\nu$ are the coefficients of the Lindblad operators. Here we suppose that each dimension holds the same coefficients $\mu$ and $\nu$ given by Eq. (13). Thus, the Quantum Langevin Dynamics in $\mathbb{R}^{d}$ space is
$\frac{\mathrm{d}\rho}{\mathrm{d}t}=-\frac{\mathrm{i}}{\hbar}[H,\rho]+\sum_{j=1}^{d}\left(2L_{j}\rho L_{j}^{\dagger}-\left[L_{j}^{\dagger}L_{j},\rho\right]_{+}\right),$ (19)
where $\rho$ is the density matrix, $H$ is the Hamiltonian given by Eq. (17), and $L_{j}$ is the Lindblad operator in Eq. (18).
3 Convergence of QLD
In this section, our primary focus is the time-independent scenario of QLD, implying that all parameters $\eta,\hbar,T,m,\Omega,k$ remain constant throughout evolution. We first discuss the convergence of QLD for the convex functions, which are defined as follows:
Definition 1.
Let $x^{*}$ be the global minimum of the differentiable function $V\colon\mathbb{R}^{d}\to\mathbb{R}$. The function $V$ is convex if for all $x\in\mathbb{R}^{d}$,
$V(x^{*})-V(x)\geq\nabla V(x)\cdot(x^{*}-x).$ (20)
In the convex case, the convergence rate of QLD is $O(e^{-\eta t})$ as demonstrated in Theorem 1. The generalization of QLD to high-dimensional convex case is quite natural while keeping the same convergent speed $O(e^{-\eta t})$, see Section 3.2. Subsequently, we compute the analytical solution of the quadratic potential, the simplest convex function, in Section 3.3. The most remarkable discovery related to the quadratic potential is that the system will reach the ground state of the quadratic potential when the evolution time $t\to+\infty$ and the temperature $T\to 0$. Lastly, we prove that QLD continues to converge to the global minimum even in a specific nonconvex scenario, where the objective function satisfies $x\cdot\nabla V(x)\geq rV(x)\geq 0\text{ for }r\in(0,1)$, at a rate of $O(e^{-\eta rt})$. More details can be referred to Theorem 3 of Section 3.4.
3.1 Convergence of QLD in the convex case
In this subsection, we mainly focus on the one-dimensional case. Our goal is to demonstrate that the expectation value of the one-dimensional objective function $V(x)$ converges to the global minimum $V(x^{*})$ at a defined rate under the evolution of QLD. This necessitates the consideration of the expectation value of the Hamiltonian $H$, as if we can establish that the expectation value of $H$ diminishes, it follows that the expectation value of $V$ will also decrease. With this understanding, we initially present Lemma 1, which enables us to calculate the time evolution of the expectation value of any observable evolving by standard Lindblad equation. After that, we show some commutation relations used in the calculation referring to Lemma 2. Subsequently, we verify the convergence of the expectation value of $H$ in Proposition 2, and with the assistance of this proposition, we validate the convergence of the expectation value of $V$ in Theorem 1.
@@ 3.1 Convergence of QLD in the convex case
$\left\langle{V}\right\rangle_{t}-V(x^{*})\leq 2kT+O(e^{-\eta t}),$ (21)
if the initial density matrix $\rho(0)$ satisfies $\left\langle{H}\right\rangle_{0}-2kT\geq 0$, where $\langle\cdot\rangle_{t}$ denotes the expectation value with respect to the density matrix $\rho(t)$, i.e., $\langle\cdot\rangle_{t}=\mathrm{Tr}(\cdot\rho(t))$.
@@
$\left\langle{V}\right\rangle_{t}\leq 2kT+O(e^{-\eta t}).$ (22)
Because $\left\langle{\frac{p^{2}}{2m}}\right\rangle_{t}+\left\langle{V}\right\rangle_{t}=\left\langle{H}\right\rangle_{t}$ and all these three terms are greater than or equal to zero, we have
$\left\langle{V}\right\rangle_{t}\leq\left\langle{H}\right\rangle_{t}.$ (23)
By Proposition 2, we get $\left\langle{V}\right\rangle_{t}\leq\left\langle{H}\right\rangle_{t}\leq 2kT+O(e^{-\eta t})$. ∎
@@
If the desired accuracy is $\varepsilon$, which means we want $\left\langle{V}\right\rangle_{t}-V(x^{*})\leq\varepsilon$ when $t\to+\infty$. To achieve this, we only need to set $kT=\varepsilon/2$ and ensure the approximation $\hbar\Omega\ll kT$ still holds.
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\langle O(t)\rangle_{t}=\left\langle{\frac{\mathrm{d}}{\mathrm{d}t}O(t)}\right\rangle_{t}+\frac{\mathrm{i}}{\hbar}\langle[H,O(t)]\rangle_{t}+\sum_{j=1}^{m}\left\langle\left(-L_{j}^{\dagger}\left[L_{j},O(t)\right]+\left[L_{j}^{\dagger},O(t)\right]L_{j}\right)\right\rangle_{t}.$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\langle O(t)\rangle_{t}$ $=\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{Tr}(O(t)\rho(t))$
$=\mathrm{Tr}\left(\frac{\mathrm{d}O(t)}{\mathrm{d}t}\rho(t)\right)+\mathrm{Tr}\left(O(t)\frac{\mathrm{d}\rho(t)}{\mathrm{d}t}\right)$
$=\left\langle{\frac{\mathrm{d}}{\mathrm{d}t}O(t)}\right\rangle_{t}+\mathrm{Tr}\left(O(t)\frac{\mathrm{d}\rho(t)}{\mathrm{d}t}\right).$ (24)
Let’s focus on the second term $\mathrm{Tr}\left((O(t)\frac{\mathrm{d}\rho(t)}{\mathrm{d}t}\right)$. Inserting Eq. (11) into the second term, we get
$\mathrm{Tr}\left(O(t)\frac{\mathrm{d}\rho(t)}{\mathrm{d}t}\right)$
$=$ $\mathrm{Tr}\left(-\frac{\mathrm{i}}{\hbar}O(t)[H,\rho(t)]+\sum_{j=1}^{m}O(t)\left(2L_{j}\rho(t)L_{j}^{\dagger}-L_{j}^{\dagger}L_{j}\rho(t)-\rho(t)L_{j}^{\dagger}L_{j}\right)\right)$
$=$ $-\frac{\mathrm{i}}{\hbar}\mathrm{Tr}\left(O(t)[H,\rho(t)]\right)+\mathrm{Tr}\left(\sum_{j=1}^{m}O(t)\left(2L_{j}\rho(t)L_{j}^{\dagger}-L_{j}^{\dagger}L_{j}\rho(t)-\rho(t)L_{j}^{\dagger}L_{j}\right)\right),$ (25)
where the first term $\mathrm{Tr}\left(O(t)[H,\rho(t)]\right)$ could be simplified as follows:
$\mathrm{Tr}\left(O(t)[H,\rho(t)]\right)$ $=\mathrm{Tr}\left(OH\rho\right)-\mathrm{Tr}\left(O\rho H\right)$
$=\mathrm{Tr}\left(OH\rho\right)-\mathrm{Tr}\left(HO\rho\right)$
$=\mathrm{Tr}\left([O,H]\rho\right)$